Please show all work, step by step, so I can understand what to do.

Ten trials are conducted in a Bernoulli process in which the probability of success in a given trial is 0.3.

If x = the number of successes, determine the following:

a. E(x) b. σx c. P(x = 5) d. P(4 ≤ x ≤ 8) e. P(x > 4)
Answer #1

This is Bernoulli process with 10 trial with probability 0.3 for each trial.

Therefore,

X ~ Bin(n,p)

Where n = 10 , p = 0.3

P(X) = nCx * pX * ( 1 - p)n-x

a)

E(X) = np

= 10 * 0.3

= 3

b)

x = sqrt [ np (1-p) ]

= sqrt [ 10 * 0.3 * 0.7 ]

= 1.4491

c)

P( X = 5) = 10C5 * 0.305 * ( 1 - 0.30)5

= 0.1029

d)

P(4 <= X <= 8) = P(X = 4) +P(X = 5) +P(X = 6) +P(X = 7) +P(X = 8)

= 10C4 * 0.304 * ( 1 - 0.30)6 +10C5 * 0.305 * ( 1 - 0.30)5 +10C6 * 0.306 * ( 1 - 0.30)4 +10C7 * 0.307 * ( 1 - 0.30)3

+10C8 * 0.308 * ( 1 - 0.30)2

= 0.3503

e)

P(X > 4) = 1 - P(X <= 4)

= 1 - [ P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4) ]

=1 - [ 10C0 * 0.300 * ( 1 - 0.30)10 +10C1 * 0.301 * ( 1 - 0.30)9 +10C2 * 0.302 * ( 1 - 0.30)8 + 10C3 * 0.303 * ( 1 - 0.30)7

+ 10C4 * 0.304 * ( 1 - 0.30)6 ]

= 0.1503